o b au nvertiie matrices in MaCR) under the Aroup cr r bu
Answer #1

Let

メ メ

to prove that H is a subgroup of the group G of all invertible matrices in M2(R) under multiplication[M2(R) means 2 by 2 matrices with real entries]

[to prove that H is a subgroup we have to prove

(i)identity belong to H

(ii) H is closed under multiplication ; that is if A\in H,B\in H\Rightarrow AB\in H

(iii) inverse exist in H]

proof

(i) the identity matrix

  01 since 1メ0

so identity element exist in H

(ii) to prove H is closed under matrix multiplication

let

:a, bメ0

B=\begin{bmatrix} c &0 \\ 0&d \end{bmatrix};c,d\neq 0

АЕН, В Е Н

AB=\begin{bmatrix} a &0 \\ 0& b \end{bmatrix}\begin{bmatrix} c &0 \\ 0& d \end{bmatrix}

  =\begin{bmatrix} ac &0 \\ 0& bd \end{bmatrix}

AB\in Hsinceac\neq 0;bd\neq 0   ;ac\in R,bd\in R

A\in H,B\in H\Rightarrow AB\in H

so H is closed under multiplication

(iii) to prove the existence of inverse in H

A=\begin{bmatrix} a &0 \\ 0& b \end{bmatrix}

det(A)=ab

A^{-1}=\frac{1}{ab}\begin{bmatrix} b &0 \\ 0& a \end{bmatrix}\in H [since 2 1a and also belongs to R]

so inverse exists for each element of H

hence H is a subgroup of G of all invertible matrices in M2(R) under multiplication